OpenStudy (anonymous):

Math Problem: (Don't know how to solve this. Please explain when showing how to solve it.)

OpenStudy (anonymous):

Evaluate\[(f(x+h)-f(x))/h\] and simplify if \[f(x) = x^2-2x\]

OpenStudy (anonymous):

this stuff is differentiation

OpenStudy (anonymous):

i forgot how to do it this method...but i normally use power rule \[\frac{ d }{ dx }x = x ^{n-1}\]

Nnesha (nnesha):

hmm first you need to find f(x+h) so substitute x for (x+h) in f(x) function \[\huge\rm f(x+h)=\color{reD}{x}^2-2\color{ReD}{x}\]

OpenStudy (anonymous):

so i need to solve for x using that equation u gave me nnesha?

OpenStudy (anonymous):

and i dont know that power rule saseal

Nnesha (nnesha):

not for x simplify that that's the equation i gave you you already have the formula

Nnesha (nnesha):

and i don't know that either :(

Nnesha (nnesha):

\[\huge\rm f(x+h)=\color{reD}{(x+h)}^2-2\color{ReD}{(x+h)}\] solve that use foil method for (x+h)^2 and distribute parentheses by -2

OpenStudy (anonymous):

ok...power rule is like (ill just show for x^2) \[\frac{ d }{ dx }(x^2) = 2x\]

OpenStudy (anonymous):

How did you get 2x? what was the value of d @saseal?

OpenStudy (anonymous):

actually they call it dy/dx it means to differentiate

OpenStudy (anonymous):

So the answer for that problem would be\[(x ^{2} +2xh +h ^{2})-2x-2h\] @Nnesha?

OpenStudy (anonymous):

so are you suppose to plug in y and x instead of d @saseal?

Nnesha (nnesha):

no not the answer that's just f(x+h) now use that formula \[\huge\rm \frac{ f(x+h)-f(x) }{ h }\] replace f(x+h) by (x^2 +2xh+h^2)-2x-2h and f(x) by the function x^2-2x

Nnesha (nnesha):

\[\huge\rm \frac{ \color{ReD}{f(x+h)}-\color{blue}{f(x)} }{ h }\] \[\huge\rm \frac{ \color{reD}{ x^2 +2xh+h^2-2x-2h } -\color{blue}{(x^2 -2x) }}{ h }\] simplify !

OpenStudy (anonymous):

\[f'(x)=\frac{((x+h)^2-2x+h)- (x^2-2x)}{ h}\]

OpenStudy (anonymous):

(2xh+h^2-2h)/h

OpenStudy (anonymous):

That's what I got after putting like terms together

OpenStudy (anonymous):

ill show you the power rule magic after you get everything together

Nnesha (nnesha):

there is a common factor h^2+2xh+h

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Oh ya. h(h+2x+1)

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Am I suppose to factor or insert it into another equation?

Nnesha (nnesha):

\[\large\rm \frac{ \color{reD}{ \cancel{x^2} +2xh+h^2\cancel{-2x}-2h } -\color{blue}{\cancel{x^2} +\cancel{2x }}}{ h }\] \[\large\rm \frac{ h^2+2xh+h }{ h}\] \[\large\rm \frac{ h(h+2x+1)}{ h}\]

Nnesha (nnesha):

can you cancel out h?

OpenStudy (anonymous):

Oh whoops I forgot about the h on the bottom. So the answer would actually be h+2x+1?

Nnesha (nnesha):

yep

OpenStudy (anonymous):

Alright thanks nnesha and saseal!

OpenStudy (anonymous):

where did the coefficient of -2h go?

Nnesha (nnesha):

hmm let's see.

OpenStudy (anonymous):

I think we wrote that step wrong.

OpenStudy (anonymous):

Alright its 2x-2+h

Nnesha (nnesha):

\[\large\rm \frac{ \color{reD}{ \cancel{x^2} +2xh+h^2\cancel{-2x}-2h } -\color{blue}{\cancel{x^2} +\cancel{2x }}}{ h }\] \[\large\rm \frac{ h^2+2xh\color{ReD}{-2h} }{ h}\] \[\large\rm \frac{ h(h+2x\color{reD}{-2})}{ h}\] i forgot it suppose to be -2h not just h

Nnesha (nnesha):

thanks saseal

Nnesha (nnesha):

\[\large\rm \frac{ \cancel{h}(h+2x\color{reD}{-2})}{\cancel{ h}}\]

OpenStudy (anonymous):

Ok, as promised the power rule: you take the power of the variable and pull it down and minus 1 to the power like this \[(\frac{ dy }{ dx })2x^2 = (2)2x ^{2-1}\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

if its just 2x it becomes \[\frac{ dy }{ dx }(2x ^{1}) = 2x^0=2(1)\]

OpenStudy (anonymous):

why did it become 2x^0?

OpenStudy (anonymous):

so for your equation x^2-2x \[\frac{ dy }{ dx }(x^2-2x) = 2x-2\]

OpenStudy (anonymous):

because x^1 is x, after you differentiate it you need to -1 from its power so it becomes x^0

OpenStudy (anonymous):

Okay I think I understand what you did. Would we learn this in AP Calculus AB?

OpenStudy (anonymous):

if you wonder where the h went...its removed by limits as h approaches 0

OpenStudy (anonymous):

Yea you would

OpenStudy (anonymous):

it's like the meat of the calculus

OpenStudy (anonymous):

Oh okay. Did you take AP Calculus?

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

No, AP is not available in my country, we do A levels instead which is about the same

OpenStudy (anonymous):

i don't live in us

OpenStudy (anonymous):

Oh okay sorry about that.